OpenStudy (anonymous):

Amber falls out of the 17th floor of a building. The window is 144 ft about a metal ventilator box. Upon impact, he crushes it to a depth of 40 in. What was his speed before impact? What was his average acceleration when colliding with the box? How long did the collision last? Did he live, blackout, or die? *Convert 144 ft to meters.

OpenStudy (anonymous):

Convert to metres, and calculate the velocity after that many metres with acceleration of 9.8m/s/s. Ignore air resistance. Calculate his change in velocity from whatever that is to 0 over 40 inches (in metres) to work out how long the average acceleration during that time and how long the collision lasted.

OpenStudy (mathmate):

"above" 144'=144*.3048m=43.9m Using g=9.81 m/s^2 (acceleration due to gravity) v1=final velocity v0=initial velocity S=distance travelled v1^2=v0^+2gS =0^2+2*9.81*43.9=861.1 v1=29.35 m/s Again using v2^2=v1^2+2aS (a=acceleration, negative for deceleration) 0=29.35^2+2a*(40") a=-29.35/(2*40") =-29.35/(2*1.016m) =-14.4 m/s^2 To find the time, use v2=v1+at t=(v2-v1)/a=(0-29.35)/(-14.4)=2.03 s. If he landed on his feet, he might survive.